OpenStudy (anonymous):

how do i graph the equation and find the y intercept of y=1/2x?

OpenStudy (anonymous):

Finding the y-intercept first usually works better. To find the y-intercept, let x = 0 and find the resulting y-value. This is also zero. This means that the y-intercept occurs where x = 0 and y = 0 (at the point (0,0)). Then plot more points by putting more x's in. Let x = 2 to get y = 1 (2, 1) is the point. Let x = 4 to get y = 2 (4, 2). Then connect the dots to form the straight line.

OpenStudy (anonymous):

i'm a little confused, so the y intercept is 0? i understand plotting the points, thanks!

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok i think i understand. so what happens in the equation y=-2/3x+5? is the y intercept 5? and how do i graph that equation?

OpenStudy (anonymous):

The y-intercept is 5. If you let x = 0, then you would just get y = 5 as your y-intercept. Graph the equation by picking more x-values, getting y-values, and connecting the dots.

OpenStudy (anonymous):

ok, i'm getting the hang of it. so for 5x-15=3y when i plug in the zeros, my first point would be (3,0) and the second would be (0,-5) then i connect them with a line.

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

great, thanks!

OpenStudy (anonymous):

if i multiply the fraction -2/3 by 2, do i get -4/3?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for y=-2x+3, would my slope be -2 since the formula is y=mx+b?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

i think i'm getting this!

OpenStudy (anonymous):

so for the equation x=5y=10, is the slope -1/5?

OpenStudy (anonymous):

x - 5y = 10... the slope is going to be a positive 1/5. You subtract x on both sides to get -5y = -x + 10. Dividing by -5 gives you y = 1/5x - 2.

OpenStudy (anonymous):

Oh, whoops. You were right. Confused on the equals sign typo. If x + 5y = 10... then 5y = -x + 10, and y = -1/5x + 2

OpenStudy (anonymous):

lol, ok great!

OpenStudy (anonymous):

so going back to the equation y=-2/3x+5, i'm a little confused doing my math to get points to plot on the graph ... i keep getting fractions when i multiply and it makes it hard to plot a point.

OpenStudy (anonymous):

If you pick multiples of 3 (like 3, 6, 9, etc.), it'll cancel out the 3 on the bottom. 3 * 2/3 for example, gives 2. -6 * 2 / 3 = -2. Now you have a -2/3, so you'll get different results when you multiply it by 3, 6, 9, etc.

OpenStudy (anonymous):

Whoops, -6 * 2 / 3 = -4

OpenStudy (anonymous):

oh ok, that's a nice shortcut. thanks